OpenStudy (anonymous):

i need to find a formula for f^-1(x) when f(x)=x^3

geerky42 (geerky42):

f(x) = y = x³ Now switch x and y to each other. x = y³ This is your inverse function. Now isolate y.

geerky42 (geerky42):

Still not clear?

geerky42 (geerky42):

@SugarRainbow

OpenStudy (anonymous):

sorry i really quick

OpenStudy (anonymous):

i was eating

OpenStudy (anonymous):

y=[\[Y=\sqrt[3]{X}\]x^(1/3)]

geerky42 (geerky42):

Yup, this is your inverse function.

OpenStudy (anonymous):

i actually got Y=X−−√3

OpenStudy (anonymous):

i mean i got what kevinbourne thied got except for the x^(1/3) how did he get that?

geerky42 (geerky42):

IT means same thing, just another way to write \(\Large \sqrt[3]{x}\)

OpenStudy (anonymous):

is ti better to write it the other way or does it not really matter?

geerky42 (geerky42):

I don't think it matters.

OpenStudy (anonymous):

okay so for this next question i need to find the inverse of f(x)=x^3+5 i got x^(1/3)-5 correct?

OpenStudy (anonymous):

x=y3+5 y3=-x-5 y=[(-x)^(1/3)]-5

geerky42 (geerky42):

I believe it's (x-5)^(1/3)

OpenStudy (anonymous):

what why -x? more explanation please from kevinbournethird and geerky tell me how you got yours?

OpenStudy (anonymous):

I put my steps in my answer.

OpenStudy (anonymous):

i know but how and why did you get -x?

geerky42 (geerky42):

Do the same thing: switch x and y to each other. x³ + 5 = y y³ + 5 = x y³ = x - 5 y = ³√(x - 5)

OpenStudy (anonymous):

but hold on let me check somethin

OpenStudy (anonymous):

wait that is what i got except i didn't group the x-5 together in parentheses why u do that?

geerky42 (geerky42):

So we know that '-5' is part of the cube root. You also can write it like that: \(\large y = \sqrt[3]{x-5}\)

geerky42 (geerky42):

x^(1/3)-5 looks like \(\sqrt[3]{x} - 5\)

OpenStudy (anonymous):

yeah but see i was thinkin of it like you find the 3root of x and then subtract five or is it a little backwards since it IS an inverse and ervrythings backwards?

geerky42 (geerky42):

well, inverse is just a function that was flipped over line y=x so we switch x and y to each other then isolate y.

OpenStudy (anonymous):

okay then so the next question is  so would it be like the previous one and be x^3-5?